For the fiscal year​ 2007, a tax authority audited 1.87​% of individual tax returns with income of​ $100,000 or more. Suppose this percentage stays the same for the current tax year. What is the probability that two randomly selected returns with income of​ $100,000 or more will be​ audited?
Answer #1

due to independence:

probability that two randomly selected returns with income of​ $100,000 or more will be​ audited

=P(first will be audited)*P(2nd will be audited) =0.0187*0.0187 =0.000350